### v1

# Market Reaction to News Flows in Supply Chain Networks

Abstract
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys. We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
Introduction
Firms interact with and influence each other through various types of networks. Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business practices is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other. [11]
Global supply chains have rapidly expanded because interconnected firms through supply chains can greatly benefit from the network. Most notably, supply chains can empower firms by enhancing efficiency of production processes and thus optimizing firms’ economic performance.[12, 13] For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity.[14] Moreover, it is often evidenced that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology.[15, 7, 16]
Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability.[17, 18, 19] Such negative shocks examined in the literature include the Great East Japan Earthquake in 2011,[20] the massive flood in Thailand in 2011,[21] the COVID-19 pandemic from 2020 to 2022,[22], and the Russia-Ukraine war that started in 2022.[23]
Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains. Because the corporate value of publicly listed firms is evaluated in the stock market, [24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock price of its suppliers and customer firms. For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its customers, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31]. Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and customers, such as Aisin.[32, 33]
These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains. Although previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], no study has investigated the diffusion of shocks to other firms’ stock prices through supply chains to the best of our knowledge. To fill the gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms. Specifically, we test whether the level of negative sentiment about a firm in a news article has a positive impact on the change rate of the stock price of the firm and its suppliers and customers before and after the disclosure of the news.
We contribute to the literature by utilizing large-scale data in many respects. First, our firm-level data covers all publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys. Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles. Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news. Another notable contribution of this study is that we utilize sentiment analysis to determine the positiveness of each news article. Prior research on sentiment analysis has predominantly focused on neural networks [37, 38], and recently, Bidirectional Encoder Representations from Transformers (BERT)[39] has garnered the most acclaim as the state-of-the-art approach [35]. We take advantage of FinBERT,[40] an extended model of BERT particularly for financial information.
Data
This study utilizes four major data sources. First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016. The information in each news article includes the date, headline, main text, International Securities Identification Numbers (ISIN) of firms mentioned in the article, and language. As we will explain later in detail, our sentiment analysis to determine whether each news article is positive or negative about firms mentioned there utilizes the main text. The average number of words in an article is 243, whereas its median is 136 and maximum is 3911. Although the total number of news articles in the NewsArchve for the period examined is 20,803,561, we focus on 3,447,425 news articles that mention any firm publicly listed in 105 stock markets in the world. 90% of these news articles mention 3 or fewer firms, whereas the maximum number of mentioned firms in an article is 33 (Figure 1). The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed. **% of firms were never mentioned, and **% were mentioned only once, while 0.1% were mentioned more than 10,000 times (Figure 2).
Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API). We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API. We particularly define the daily stock price of each firm as its closing price. In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend. The Refinitiv index is available for 51 major markets that cover 92% of listed firms worldwide.
Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003-2016 obtained as of January 18, 2023. The FactSet data include major clients of each firm reported in its financial statement. The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [41]. Therefore, based on published financial statements and, in addition, supplemental information from news articles and web sites, FactSet identifies supply chains of listed and non-listed firms with their major partners. Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year. The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed more recently and thus ignores firms in the distant past.
| 2003 | 2004 | 2005 | 2006 | 2007 | 2008 | 2009 | |
| Firms across the world (total) | |||||||
| Number of firms | 9,283 | 9,292 | 9,804 | 10,058 | 10,342 | 10,197 | 10,802 |
| Number of links | 38,594 | 40,749 | 45,172 | 38,866 | 34,894 | 33,136 | 28,640 |
| firms across the world (listed) | |||||||
| Number of firms | 633 | 726 | 783 | 797 | 745 | 760 | 1,047 |
| Number of links | 1,413 | 1,685 | 1,914 | 1,723 | 1,386 | 1,364 | 1,687 |
| 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
| Firms across the world (total) | |||||||
| Number of firms | 17,729 | 24,671 | 28,244 | 31,090 | 40,146 | 53,114 | 64,968 |
| Number of links | 42,971 | 61,675 | 72,671 | 82,673 | 101,748 | 133,134 | 167,772 |
| firms across the world (listed) | |||||||
| Number of firms | 2,153 | 3,842 | 4,798 | 5,779 | 8,418 | 10,729 | 12,728 |
| Number of links | 4,179 | 8,782 | 12,073 | 15,485 | 22,473 | 31,537 | 42,201 |
Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008-2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan. Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm. Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses. However, these links can be mostly identified by responses by their suppliers and clients. Accordingly, the maximum number of suppliers and clients is *** and ***, respectively, in 2016, implying that our data cover supply chains of Japanese firms to a great extent. The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
| 2008 | 2009 | 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
| Japan (total) | |||||||||
| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
| Japan (listed firms) | |||||||||
| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 | 3,215 | 3,204 | 3,215 | 3,205 |
| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 | 21,696 | 21,439 | 20,968 | 20,445 |
Methods
Sentiment analysis using FinBERT
Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[40]. FinBERT is an extended version of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al. (2018)[39]. Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database. From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the probabilities of the three about the firms mentioned in the article. Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end.
The distribution of the probabilities of positive, neutral, and negative sentiments generated by FinBERT from news articles about firms across the world and Japanese firms are provided in Panels (A) and (B) of Figure 3, respectively. Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
From the probabilities of the three types of sentiment, we construct an index of positive sentiment by the ratio of the probability of positive sentiment to the sum of the probabilities of positive and negative sentiment, ignoring the probability of neutral sentiment. For example, when the probabilities of positive, neutral, and negative sentiment are 0.3, 0.5, and 0.2, respectively, our positive sentiment index is . We ignore the probability of neutral sentiment to avoid the ambiguity of sentiment in news articles. Figure 4 demonstrates the distribution of this positive sentiment index for firms across the world and Japanese firms. Because we avoid the ambiguity of the sentiment, the distribution of the index is skewed to both lower (0) and upper (1) limits although we observe a hike around the value of 0.25. In other words, our positive sentiment index distinguishes between positive and negative sentiment more clearly than the probabilities of the three directly generated by FinBERT and highlights positive sentiment in news articles.
Regression analysis
Direct effect
To estimate the effect of positive sentiment in news articles about particular firms on their own stock prices, we apply the positive sentiment index constructed above to regression analysis. Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness of the article after the article is disclosed, compared with its change rate before the disclosure. Suppose that firm listed in stock market is mentioned by news article on day . Setting the time window at , such as 2, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
where is the average stock price between day and . For example, is the average stock price of firm for days between day and . Therefore, the left-hand side of equations (1) and (2) indicates the daily percentage change of the average stock price in time window in the pre- and post-news period, respectively. We take the average of stock prices over the time window to reduce their fluctuations. We exclude Saturdays and Sundays from our daily observations because stock prices are not available on weekends. Accordingly, for example, when news article about firm is disclosed on Thursday, the 10th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average on 8th (Tuesday)-10th (Thursday), as shown in Figure 5. Similarly, the post-news change is defined as the change from the average on June 7th-9th to that on June 10th, 11th (Friday), and 14th (Monday).
Thus, our sample is restricted to firm-day observations before or after news articles about the firms are disclosed. In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2. Using the pre- and post-news observations, we estimate the following estimation equation:
where and are indicator variables that take a value of one if is the percentage change of the stock price of firm before and after news article is disclosed on day , respectively, or . represents the level of positive sentiment in news article generated by FinBERT (Data section). is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies. Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industry Classification at the two-digit level for the sample of Japanese firms.
Accordingly, indicates the average effect of the positiveness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors. By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through other informal channels before the disclosure of the news. In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned. Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [42].
We particularly estimate equation (3) for to examine both the short- and long-run effect of positive news articles on stock prices. We hypothesize that a positive news article would accelerate the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run. Note that equation (3) for cannot be estimated because the change rates of the stock price in the pre- and post-news periods are the same, according to equations (1) and (2).
It should be emphasized that this analysis is comparing firms positively mentioned in news articles with those negatively mentioned, rather than comparing firms positively mentioned and those not mentioned by news articles. We focus on the sub-sample of firms mentioned by news articles and ignore those not mentioned, because the two types of firms are most likely to be intrinsically different from each other to a great extent so that comparing the two may be associated with large endogeneity biases.
Indirect effect through supply chains
We further examine the effect of positive sentiment about firms in news articles on their suppliers and customers, using similar frameworks. Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive sentiment in the news article that mentions not firm but any of its customer firms. Alternatively, when we examine the downstream effect of customers’ news on their customers, is replace with the positive index of any of suppliers of firm . When a firm mentioned by a news article has more than one supplier or customer, as shown in Tables 1 and 2, we include all the suppliers and customers in our sample. In this analysis, we compare stock prices of suppliers or customers of firms mentioned positively in news articles with those of supplier or customers of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
We hypothesize that the effect of news articles on stock prices diffuses through supply chains. Precisely, a positive news article about a customer firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the customer firm. Also, positive news about a supplier would promote growth of its customers because of possible productivity growth of the supplier. As in the case of the direct effect, the indirect effect on suppliers and customers would be smaller in the long run than in the short run.
Results and Discussion
Direct effect of news articles on listed firms across the world
We start with estimating how the positiveness of news articles about firms affect percentage changes in their stock prices before and after the disclosure of news, applying equation (3) to the sample of publicly listed firms across the world mentioned by any news article. Results are presented in Table 3 and Figure 6.
| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | ||||||
| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) |
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Number of observations | 9,286,677 | 9,334,585 | 9,354,107 | 9,369,523 | 9,402,470 | 9,381,512 | 9,314,442 |
In column (1) of Table 3, the estimated is 0.419 and highly significant, indicating that the level of the positiveness in a news article about a firm is positively correlated with the percentage change in the firm’s stock price for 2 days before the disclosure of the news. Specifically, because the positiveness index, in equation (3), ranges from 0 to 1 (Figure 4), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.419% points higher than that of a firm mentioned most negatively () before the disclosure of the news. There are two possibilities for this positive coefficient. First, positive and negative information included in news articles diffuses to stock markets partially through informal channels before the formal disclosure of news, generating the pre-news effect. Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms negatively mentioned, leading to endogeneity biases.
Column (1) of Table 3 further finds that the estimated is 0.505, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.505% points than that of a firm mentioned most negatively after the news disclosure. The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0. In column (1), the value is close to 0, and thus we reject the null hypothesis. This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned negatively even before the disclosure of news, positive news raises stock prices of former firms more after its disclosure.
These results in column (1), or for the time window of 2 days, generally hold in columns (2)-(5) for time windows of 3-5 and 30 days, while the post- and pre-news effects and the difference between the two are smaller as the time window becomes wider. By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size. This result implies that the intrinsic difference between firms mentioned positively and negatively by news articles is negligible. Therefore, the positive and significant difference between the post- and pre-news effects in the short run (time windoes from 2 to 30 days) is mostly due to information diffusion through informal channels before the disclosure of news.
Figure 6 graphically demonstrates the argument above. The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect. Although the confidence intervals are added in the figure, they are too small to be recognized. The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges. This shape suggests the followings. The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news. However, the change rate of the stock price of positively-mentioned firms started to increase compared with negatively-mentioned firms 30 days before the disclosure and reaches the peak 2 days after the disclosure. Then, the difference between the change rates of the stock prices of positively- and negatively-mentioned firms becomes smaller over time and converges to almost 0 180 days after the disclosure.
Indirect effect on suppliers through global supply chains
Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and customers through global supply chains by estimating equation 3 where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or customer of firm .
Figure 7 is the graphical presentation of the results for the effect of news about firms on their suppliers, whereas Table 4 particularly shows the difference between the post- and pre-news effects and its values. The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 6 and Table 3). The post- and pre-news effects are positive and significant, and their difference is positive and significant when the time window is narrow. These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news through informal information diffusion and raises them more after the disclosure.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| value | 0.00 | 0.00 | 0.23 | 0.00 | 0.00 | 0.03 | 0.00 |
However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers. First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference. For example, the pre-news effect on suppliers for the time window of 2 days is 0.0238 (Figure 7), while the corresponding direct effect is 0.419 (Figure 6), more than 10 times larger. Similarly, the net effect of the disclosure of news, represented by and shown in Table 4, is 0.0033 when the time window is 2 days, while the corresponding value for the direct effect is 0.086 (Table 3). Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms. Second, the difference between the post- and pre-news effects on suppliers is positive and significant only when the time window is 2 or 3 days, although the significantly positive difference between the post- and pre-news direct effects is observed in a longer run, up to the time window of 30 days. This implies that the net effect of the disclosure of news on suppliers lasts for only several days, a very short run, after the disclosure.
Indirect effect on customers through global supply chains
Similarly, we examine the effect on customer firms and show the results in Figure 8 and Table 5. In this analysis, although the pre- and post-news effects on customers show an inverted-U shape (panel [A] of Figure 8), as in the previous results, we do not find any positive and significant difference between the post- and pre-news effects (the upper part of Table 5). Therefore, in addition to the benchmark index of positiveness of each news article using the probabilities of positive and negative sentiment from the FinBERT (Figure 4), we employ an alternative positiveness index based on only the probability of positive sentiment shown by pink bars in Figure 3. As is clear from the comparison between Figures 3 and 4, the alternative index tends to evaluate the positiveness of news articles less than the benchmark index. Thus, using the alternative index, we can examine the effect of highly positive news. Then, we find a positive and significant difference between the post- and pre-news effects, or the net effect of the disclosure of highly positive news articles, on customers for the time windows of 2 and 3 days, as shown in the lower rows of Table 5. As in the case of the effect on suppliers, the net effect on customers is small and lasts only for several days.
| 2 | 3 | 4 | 5 | 30 | 180 | 365 | |
| Benchmark positiveness index using the positive and negative probabilities of news articles | |||||||
| 0.0008 | -0.0007 | ||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Alternative positiveness index using only the positive probability of news articles | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
Effect on Japanese firms
We further use another sample that focuses on Japanese listed firms. The results for the direct effect of positive news articles on Japanese firms’ own stock prices shown in Figure 9 are qualitatively the same as and quantitatively similar to those using the sample of firms across the world shown in Figure 6.
In addition, we examine the indirect effect of news articles about firms on stock prices of their suppliers and customers. A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section). Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
The results are presented in Figure 10 and Table 6. Panels (A) and (B) of Figure 10 for the effect on suppliers and customers, respectively, show several characteristics similar to each other but different from the previous results. First, although the pre- and post-news effect on stock prices of both suppliers and customers are mostly positive and significantly different from 0, as in the previous results, the pre-news effect is negative when the time window is 180 and 365 days, unlike the previous results. The negative pre-news effect in the long run implies that stock prices of firms positively mentioned in news do not necessarily grow faster long time before the news. In other words, the effect of news is less likely to be overestimated due to selection of potentially fast-growing firms.
| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
| Effect on suppliers | |||||||
| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
| Effect on customers | |||||||
| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
Second, the pre-news effect is positive and significant, implying diffusion of information before the disclosure of news that is consistent with the previous results. However, the difference between the post- and pre-news effects that can be interpreted as the net effect of the disclosure of news is negative in the short run, or in the time window of 2-5 days and turns to be positive in the long run (Table 6). This result is completely opposite to the previous ones using data for firms across countries (Tables 4 and 5). In other words, positive information about a firm diffuses through informal channels and raises the change rate of stock prices of the firm’s suppliers and customers before the information is disclosed in a news article. However, the disclosure of the information by a news article lowers the change rate of stock prices compared with that in the pre-news period, although the post-news effect is still positive. One possible reason for this is that investors who can receive information about firms through informal channels are more knowledgeable and recognize diffusion of performance through supply chains more accurately than other investors who rely on formal news. Therefore, once the knowledgeable investors receive positive information about a firm before the disclosure by news, they invest in the firm’s suppliers and customers. However, other investors react less to positive news about firms’ suppliers and customers.
Finally, the pre- and post-news effects on suppliers and clients found in the sample of Japanese firms are larger than those in the sample of firms across the world. This finding, combined with the previous interpretation about knowledgeable investors, suggests that supply chain ties between suppliers and customers in Japan, known as keiretsu, are often stronger than in other countries, because they are often associated with shareholding relationships, information exchanges, technical assistance and research collaboration.[43, 44] Therefore, diffusion of performance through supply chains may be more prominent in Japan than in others, and this possibility is recognized by at least knowledgeable investors in Japan.
Conclusion
This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms. The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information. Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels. Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains. In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect. The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion. However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world. This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
Acknowledgements
The authors thank Yuito Nishi for his technical contribution.
Author contributions statement
H.I conceived the study. H.I and Y.T conducted the analyses and wrote and reviewed the manuscript.

### v3

@@
 Abstract
-This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms.
-The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
+This study examines how positive and negative news about firms affects their stock prices and, moreover, how it affects stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms around the world and another of Japanese listed firms.
+The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
 Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
-We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels.
-Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains.
-In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect.
-The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion.
-However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world.
-This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
-Introduction
+We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through unofficial information channels.
+Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains.
+In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect.
+The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels.
+However, the post-news effect on suppliers and clients in Japan is smaller than the pre-news effect, which is the opposite result to non-domestic firms from around the world.
+keywords:
+Supply Chain, Natural Language Processing, Stock Market, Large -scale Data1 Introduction
 Firms interact with and influence each other through various types of networks.
-Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business practices is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other.
+Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other.
 [11]
@@
 Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains.
-Because the corporate value of publicly listed firms is evaluated in the stock market, [24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock price of its suppliers and customer firms.
-For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its customers, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31].
-Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and customers, such as Aisin.[32, 33]
+Because the corporate value of publicly listed firms is evaluated in the stock market,[24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms.
+For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its clients, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31].
+Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and clients, such as Aisin.[32, 33]
 These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains.
-Although previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], no study has investigated the diffusion of shocks to other firms’ stock prices through supply chains to the best of our knowledge.
-To fill the gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms.
-Specifically, we test whether the level of negative sentiment about a firm in a news article has a positive impact on the change rate of the stock price of the firm and its suppliers and customers before and after the disclosure of the news.
+Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], there are not so many studies that investigated the diffusion of shocks to other firms’ stock prices through supply chains.
+Kim and Wagner [37] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients.
+Pandit et al.
+utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyzed their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT[38].
+In addition, similar studies focusing on China are discussed [39, 40].
+However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
+To fill this gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms.
+Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
 We contribute to the literature by utilizing large-scale data in many respects.
-First, our firm-level data covers all publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys.
+First, our firm-level data covers most publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys.
 Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles.
 Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news.
-Another notable contribution of this study is that we utilize sentiment analysis to determine the positiveness of each news article.
-Prior research on sentiment analysis has predominantly focused on neural networks [37, 38], and recently, Bidirectional Encoder Representations from Transformers (BERT)[39] has garnered the most acclaim as the state-of-the-art approach [35].
-We take advantage of FinBERT,[40] an extended model of BERT particularly for financial information.
-Data
+Although we have mainly focused on how the shock affects stock prices through supply chains, it is evident that this study also relates to business applications and researches that predict changes of stock prices based on the news.
+In this field, neural networks are predominantly used.
+For example, hedge funds, such as Bridgewater Associates utilizes the neural networks for predictions[41].
+In academia, the transformer[42] is widely used[43, 44].
+In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory[45, 46, 47, 48, 49, 50, 51, 52], convolutional neural networks[53], or other deep learning methods[54, 55, 56, 57].
+We use sentiment analysis to determine the positiveness and negativeness of each news article[58, 59] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT)[60] has garnered the most acclaim[35].
+We take advantage of FinBERT[61], a fine-tuned model of BERT particularly for financial information.
+2 Data
 This study utilizes four major data sources.
 First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016.
-The information in each news article includes the date, headline, main text, International Securities Identification Numbers (ISIN) of firms mentioned in the article, and language.
-As we will explain later in detail, our sentiment analysis to determine whether each news article is positive or negative about firms mentioned there utilizes the main text.
-The average number of words in an article is 243, whereas its median is 136 and maximum is 3911.
-Although the total number of news articles in the NewsArchve for the period examined is 20,803,561, we focus on 3,447,425 news articles that mention any firm publicly listed in 105 stock markets in the world.
-90% of these news articles mention 3 or fewer firms, whereas the maximum number of mentioned firms in an article is 33 (Figure 1).
+The information in each news article includes the time, headline, main text, Refinitiv Identification Code (a ticker-like code) of firms mentioned in the article, language, etc.
+As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text.
+The average number of words in an article is 243, whereas its median is 136 and maximum is 3,911.
+Although the total number of news articles in the NewsArchive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world.
+Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33.
+(Figure 1).
 The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed.
-**% of firms were never mentioned, and **% were mentioned only once, while 0.1% were mentioned more than 10,000 times (Figure 2).
+55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
 Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API).
@@
 In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend.
-The Refinitiv index is available for 51 major markets that cover 92% of listed firms worldwide.
-Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003-2016 obtained as of January 18, 2023.
+The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
+Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023.
 The FactSet data include major clients of each firm reported in its financial statement.
@@
 Table 1 shows the number of all firms (for reference) and listed firms (our sample) in the FactSet data and their supply-chain links by year.
-The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed more recently and thus ignores firms in the distant past.
+The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past.
+In a later section, the recorded supply chain links at the time are considered to estimate the impact of the news on supply chains.
 | 2003 | 2004 | 2005 | 2006 | 2007 | 2008 | 2009 | |
-| Firms across the world (total) | |||||||
+| All firms | |||||||
 | Number of firms | 9,283 | 9,292 | 9,804 | 10,058 | 10,342 | 10,197 | 10,802 |
 | Number of links | 38,594 | 40,749 | 45,172 | 38,866 | 34,894 | 33,136 | 28,640 |
-| firms across the world (listed) | |||||||
+| Maximum indegree | 283 | 272 | 272 | 421 | 343 | 291 | 195 |
+| Maximum outdegree | 255 | 266 | 249 | 203 | 159 | 180 | 156 |
+| Listed | |||||||
 | Number of firms | 633 | 726 | 783 | 797 | 745 | 760 | 1,047 |
 | Number of links | 1,413 | 1,685 | 1,914 | 1,723 | 1,386 | 1,364 | 1,687 |
+| Maximum indegree | 31 | 36 | 30 | 24 | 22 | 19 | 17 |
+| Maximum outdegree | 35 | 41 | 47 | 56 | 43 | 44 | 35 |
 | 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
-| Firms across the world (total) | |||||||
+| All firms | |||||||
 | Number of firms | 17,729 | 24,671 | 28,244 | 31,090 | 40,146 | 53,114 | 64,968 |
 | Number of links | 42,971 | 61,675 | 72,671 | 82,673 | 101,748 | 133,134 | 167,772 |
-| firms across the world (listed) | |||||||
+| Maximum indegree | 463 | 600 | 577 | 496 | 423 | 586 | 562 |
+| Maximum outdegree | 162 | 201 | 206 | 198 | 256 | 294 | 378 |
+| Listed | |||||||
 | Number of firms | 2,153 | 3,842 | 4,798 | 5,779 | 8,418 | 10,729 | 12,728 |
 | Number of links | 4,179 | 8,782 | 12,073 | 15,485 | 22,473 | 31,537 | 42,201 |
-Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008-2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan.
+| Maximum indegree | 27 | 70 | 115 | 127 | 161 | 225 | 277 |
+| Maximum outdegree | 40 | 54 | 79 | 79 | 75 | 109 | 140 |
+Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan.
 Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm.
@@
 However, these links can be mostly identified by responses by their suppliers and clients.
-Accordingly, the maximum number of suppliers and clients is *** and ***, respectively, in 2016, implying that our data cover supply chains of Japanese firms to a great extent.
+Since those firms are relatively large, their partners report them as important suppliers or clients.
 The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
-| 2008 | 2009 | 2010 | 2011 | 2012 | 2013 | 2014 | 2015 | 2016 | |
-| Japan (total) | |||||||||
-| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
-| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
-| Japan (listed firms) | |||||||||
-| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 | 3,215 | 3,204 | 3,215 | 3,205 |
-| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 | 21,696 | 21,439 | 20,968 | 20,445 |
-Methods
+| 2008 | 2009 | 2010 | 2011 | 2012 | |
+| All firms | |||||
+| Number of firms | 1,018,570 | 1,061,051 | 1,107,760 | 1,135,200 | 1,156,205 |
+| Number of links | 4,443,677 | 4,644,977 | 4,810,967 | 4,925,307 | 4,995,612 |
+| Maximum indegree | 6,751 | 6,976 | 7,069 | 7,332 | 8,398 |
+| Maximum outdegree | 11,140 | 11,150 | 11,155 | 11,201 | 11,287 |
+| Listed | |||||
+| Number of firms | 3,554 | 3,456 | 3,403 | 3,308 | 3,245 |
+| Number of links | 21,720 | 22,128 | 22,630 | 22,044 | 21,833 |
+| Maximum indegree | 201 | 189 | 180 | 162 | 151 |
+| Maximum outdegree | 305 | 277 | 259 | 243 | 225 |
+| 2013 | 2014 | 2015 | 2016 | |
+| All firms | ||||
+| Number of firms | 1,182,729 | 1,194,034 | 1,200,162 | 1,247,939 |
+| Number of links | 5,090,119 | 5,157,181 | 5,209,604 | 5,491,417 |
+| Maximum indegree | 8,787 | 9,520 | 9,948 | 12,016 |
+| Maximum outdegree | 11,490 | 11,586 | 11,588 | 12,729 |
+| Listed | ||||
+| Number of firms | 3,215 | 3,204 | 3,215 | 3,205 |
+| Number of links | 21,696 | 21,439 | 20,968 | 20,445 |
+| Maximum indegree | 148 | 145 | 143 | 141 |
+| Maximum outdegree | 216 | 201 | 195 | 190 |
+3 Methods
 Sentiment analysis using FinBERT
 Our sentiment analysis to determine the degree of positiveness and negativeness of the information about the listed firms provided in each news article utilizes a natural language processing (NLP) model called FinBERT[61].
-FinBERT is an extended version of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al.
+FinBERT is a variant of the Bidirectional Encoder Representations from Transformers (BERT) developed by Devlin et al.
 (2018)[60].
 Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database.
-From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the probabilities of the three about the firms mentioned in the article.
+From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three about the firms mentioned in the article.
+Examples are shown in Table 3.
 Before conducting the sentiment analysis using FinBERT, we clean news articles by deleting URLs, line breaks, ISINs, and fixed phrases at the end.
@@
 Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
-From the probabilities of the three types of sentiment, we construct an index of positive sentiment by the ratio of the probability of positive sentiment to the sum of the probabilities of positive and negative sentiment, ignoring the probability of neutral sentiment.
-For example, when the probabilities of positive, neutral, and negative sentiment are 0.3, 0.5, and 0.2, respectively, our positive sentiment index is .
-We ignore the probability of neutral sentiment to avoid the ambiguity of sentiment in news articles.
-Figure 4 demonstrates the distribution of this positive sentiment index for firms across the world and Japanese firms.
-Because we avoid the ambiguity of the sentiment, the distribution of the index is skewed to both lower (0) and upper (1) limits although we observe a hike around the value of 0.25.
-In other words, our positive sentiment index distinguishes between positive and negative sentiment more clearly than the probabilities of the three directly generated by FinBERT and highlights positive sentiment in news articles.
-Regression analysis
+| Sentiment | |||
+| News article | Positive | Neutral | Negative |
+| (Sep/11/2008) Advanced Medical Solutions Group Plc on Thursday said its silver anti-microbial wound gel had been approved by the U.S.
+Food & Drug Administration (FDA), sending its shares up 5 percent.
+… [Source: [63]] | 0.93 | 0.01 | 0.06 |
+| (Sep/11/2008) U.S. Interior Department employees who oversaw oil drilling on federal lands had sex and used illegal drugs with workers at energy companies … [Source: [64]] | 0.02 | 0.21 | 0.76 |
+| (Sep/11/2008) Amazon.com, the largest global online retailer, plans to start selling U.S.-produced wine on its website within the United States by late September or October … [Source: [65]] | 0.14 | 0.84 | 0.01 |
+3.1 Regression analysis
 Direct effect
-To estimate the effect of positive sentiment in news articles about particular firms on their own stock prices, we apply the positive sentiment index constructed above to regression analysis.
-Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness of the article after the article is disclosed, compared with its change rate before the disclosure.
+To estimate the effect of positive and negative sentiment in news articles about particular firms on their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis.
+Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure.
 Suppose that firm listed in stock market is mentioned by news article on day .
-Setting the time window at , such as 2, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
-where is the average stock price between day and .
+Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
+where is the average stock price of firm between day and .
 For example, is the average stock price of firm for days between day and .
@@
 We take the average of stock prices over the time window to reduce their fluctuations.
-We exclude Saturdays and Sundays from our daily observations because stock prices are not available on weekends.
-Accordingly, for example, when news article about firm is disclosed on Thursday, the 10th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average on 8th (Tuesday)-10th (Thursday), as shown in Figure 5.
-Similarly, the post-news change is defined as the change from the average on June 7th-9th to that on June 10th, 11th (Friday), and 14th (Monday).
-Thus, our sample is restricted to firm-day observations before or after news articles about the firms are disclosed.
+We skip days when stock prices are unavailable, such as Saturdays, Sundays, and holidays.
+Accordingly, for example, as shown in Figure 4, when news article about firm is disclosed on Friday, the 11th of June and the time window is set at 3 days, the pre-news change rate is defined as the change rate from the average stock price on June 3rd (Thursday), 4th (Friday), and 7th (Monday), skipping 5th (Saturday) and 6th (Sunday), to the average from the 8th (Tuesday) to 10th (Thursday).
+Similarly, the post-news change is defined as the change from the average of June 8–10 to that of June 11th, 14th (Monday), and 15th (Tuesday).
+Thus, our sample is restricted to firm-day observations that occur before or after news articles about the firms are disclosed.
 In other words, the number of observations in our sample is the cumulative number of firms mentioned in news articles during the period examined multiplied by 2.
 Using the pre- and post-news observations, we estimate the following estimation equation:
-where and are indicator variables that take a value of one if is the percentage change of the stock price of firm before and after news article is disclosed on day , respectively, or .
-represents the level of positive sentiment in news article generated by FinBERT (Data section).
+where T is either pre or post, and and are indicator variables that take a value of one if represents the percentage change in the stock price of firm before or after news article is disclosed on day , respectively.
+represents the level of either positive or negative sentiment in news article generated by FinBERT (Data section).
 is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies.
-Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industry Classification at the two-digit level for the sample of Japanese firms.
-Accordingly, indicates the average effect of the positiveness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors.
-By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through other informal channels before the disclosure of the news.
+Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
+Accordingly, indicates the average effect of the positiveness or negativeness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors.
+By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through unofficial information channels before the disclosure of the news.
+Such unofficial information channels include suggestions by financial companies to their clients, posts in Social Networking Services (SNS) by influencers, and word of mouth among investors.
 In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned.
 Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [66].
-We particularly estimate equation (3) for to examine both the short- and long-run effect of positive news articles on stock prices.
-We hypothesize that a positive news article would accelerate the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run.
-Note that equation (3) for cannot be estimated because the change rates of the stock price in the pre- and post-news periods are the same, according to equations (1) and (2).
+We particularly estimate equation (3) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run effect of positive news articles on stock prices.
+We hypothesize that a positive (negative) news article would accelerate (deteriorate) the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run.
 It should be emphasized that this analysis is comparing firms positively mentioned in news articles with those negatively mentioned, rather than comparing firms positively mentioned and those not mentioned by news articles.
@@
 Indirect effect through supply chains
-We further examine the effect of positive sentiment about firms in news articles on their suppliers and customers, using similar frameworks.
-Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive sentiment in the news article that mentions not firm but any of its customer firms.
-Alternatively, when we examine the downstream effect of customers’ news on their customers, is replace with the positive index of any of suppliers of firm .
-When a firm mentioned by a news article has more than one supplier or customer, as shown in Tables 1 and 2, we include all the suppliers and customers in our sample.
-In this analysis, we compare stock prices of suppliers or customers of firms mentioned positively in news articles with those of supplier or customers of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
+We further examine the effect of positive or negative sentiment about firms in news articles on their suppliers and clients, using similar frameworks.
+Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive or negative sentiment in the news article that mentions not firm but any of ’s client firms.
+Alternatively, when we examine the downstream effect of clients’ news on their clients, is replaced with the positive or negative index of any of the suppliers of firm .
+When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample.
+In this analysis, we compare stock prices of suppliers or clients of firms mentioned positively in news articles with those of suppliers or clients of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
 We hypothesize that the effect of news articles on stock prices diffuses through supply chains.
-Precisely, a positive news article about a customer firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the customer firm.
-Also, positive news about a supplier would promote growth of its customers because of possible productivity growth of the supplier.
-As in the case of the direct effect, the indirect effect on suppliers and customers would be smaller in the long run than in the short run.
-Results and Discussion
+Precisely, a positive news article about a client firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the client firm.
+Also, positive news about a supplier would promote growth of its clients because of possible productivity growth of the supplier.
+As in the case of the direct effect, the indirect effect on suppliers and clients would be smaller in the long run than in the short run.
+4 Results and Discussion
 Direct effect of news articles on listed firms across the world
@@
 Results are presented in Table 4 and Figure 5.
-| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | ||||||
-| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) |
-| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
-| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
-| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
-| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
-| Number of observations | 9,286,677 | 9,334,585 | 9,354,107 | 9,369,523 | 9,402,470 | 9,381,512 | 9,314,442 |
-In column (1) of Table 3, the estimated is 0.419 and highly significant, indicating that the level of the positiveness in a news article about a firm is positively correlated with the percentage change in the firm’s stock price for 2 days before the disclosure of the news.
-Specifically, because the positiveness index, in equation (3), ranges from 0 to 1 (Figure 4), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.419% points higher than that of a firm mentioned most negatively () before the disclosure of the news.
+| Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | |||||||
+| Model: | (1) | (2) | (3) | (4) | (5) | (6) | (7) | (8) |
+| Time window (days) [] | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
+| Market-specific trend | Yes | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
+| Sector fixed effects | Yes | Yes | Yes | Yes | Yes | Yes | Yes | Yes |
+| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
+| Number of observations | 9,409,978 | 9,543,586 | 9,562,439 | 9,565,054 | 9,562,897 | 9,460,176 | 8,986,209 | 8,546,395 |
+In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that the probability of positive sentiment in a news article about a firm is positively correlated with the percentage change in the firm’s stock price from 2 days to 1 day before the disclosure of the news.
+Since the positiveness measure, in equation (3), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.323% points higher than that of a firm mentioned with no positive sentiment () before the disclosure of the news.
 There are two possibilities for this positive coefficient.
-First, positive and negative information included in news articles diffuses to stock markets partially through informal channels before the formal disclosure of news, generating the pre-news effect.
-Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms negatively mentioned, leading to endogeneity biases.
-Column (1) of Table 3 further finds that the estimated is 0.505, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.505% points than that of a firm mentioned most negatively after the news disclosure.
+First, positive information included in news articles diffuses to stock markets partially through private channels before the formal disclosure of news, generating the pre-news effect.
+Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms with no positive mention, leading to endogeneity biases.
+Column (1) of Table 4 further finds that the estimated is 0.923, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.923% points than that of a firm mentioned with no positive sentiment after the news disclosure.
 The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0.
 In column (1), the value is close to 0, and thus we reject the null hypothesis.
-This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned negatively even before the disclosure of news, positive news raises stock prices of former firms more after its disclosure.
-These results in column (1), or for the time window of 2 days, generally hold in columns (2)-(5) for time windows of 3-5 and 30 days, while the post- and pre-news effects and the difference between the two are smaller as the time window becomes wider.
+This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, positive news raises stock prices of former firms more by 0.6% points after its disclosure.
+These results in column (1) for the time window of 1 day generally hold in columns (2)-(5) for time windows of 2-5 and 30 days, while the post- and pre-news effects and the difference between the two decrease as the time window becomes wider.
 By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size.
-This result implies that the intrinsic difference between firms mentioned positively and negatively by news articles is negligible.
-Therefore, the positive and significant difference between the post- and pre-news effects in the short run (time windoes from 2 to 30 days) is mostly due to information diffusion through informal channels before the disclosure of news.
-Figure 6 graphically demonstrates the argument above.
+This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
+Panel (A) of Figure 5 graphically demonstrates the argument above.
 The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect.
@@
 The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news.
-However, the change rate of the stock price of positively-mentioned firms started to increase compared with negatively-mentioned firms 30 days before the disclosure and reaches the peak 2 days after the disclosure.
-Then, the difference between the change rates of the stock prices of positively- and negatively-mentioned firms becomes smaller over time and converges to almost 0 180 days after the disclosure.
-Indirect effect on suppliers through global supply chains
-Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and customers through global supply chains by estimating equation 3 where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or customer of firm .
-Figure 7 is the graphical presentation of the results for the effect of news about firms on their suppliers, whereas Table 4 particularly shows the difference between the post- and pre-news effects and its values.
+However, the change rate of the stock price of positively mentioned firms began to rise compared to firms with no positive mention 30 days before the disclosure, peaking one day after the disclosure.
+Then, the difference in the change rates of stock prices between positively mentioned firms and those with no positive mention gradually decreases over time, converging to nearly zero 180 days after the disclosure.
+The effect of the probability of negative sentiment in each news article shown in Panel (B) of Figure 5 is consistent with the effect of the probability of positive sentiment in Panel (A).
+The effect of negative sentiment in a news article is negligible 180 days or more before its disclosure but becomes negative and significant 30 days or fewer before the disclosure, implying information diffusion through private channels.
+After the disclosure, the negative effect becomes larger in the absolute term, suggesting a negative net effect of the disclosure of negative news on stock prices, in addition to the negative effect through private channels.
+Indirect effect on suppliers and clients through global supply chains
+Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and clients through global supply chains by estimating equation (3) where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
+Panels (A) and (B) of Figure 6 are the graphical presentation of the results for the effect of positive and negative news about firms, respectively, on their suppliers.
+Table 5 particularly shows the difference between the post- and pre-news effects of positive news and its values.
+The corresponding table for the effect of negative news is omitted here but is available in the Supplementary Information (SI Table 1).
 The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 5 and Table 4).
-The post- and pre-news effects are positive and significant, and their difference is positive and significant when the time window is narrow.
-These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news through informal information diffusion and raises them more after the disclosure.
-| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
-| value | 0.00 | 0.00 | 0.23 | 0.00 | 0.00 | 0.03 | 0.00 |
+The post- and pre-news effects of positive news are positive and significant, and their difference is positive and significant when the time window is narrow.
+These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news due to information diffusion through private channels and raises them more after the disclosure due to the additional effect of news medias.
+The effect of negative news articles is opposite in the direction and similar in size to the effect of positive ones, indicating the same conclusion for negative news.
+| Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
+| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.27 | 0.00 | 0.00 |
 However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers.
 First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference.
-For example, the pre-news effect on suppliers for the time window of 2 days is 0.0238 (Figure 7), while the corresponding direct effect is 0.419 (Figure 6), more than 10 times larger.
-Similarly, the net effect of the disclosure of news, represented by and shown in Table 4, is 0.0033 when the time window is 2 days, while the corresponding value for the direct effect is 0.086 (Table 3).
+For example, the pre-news effect on suppliers for the time window of 1 days is 0.0293 (Figure 6), while the corresponding direct effect is 0.323 (Figure 5), more than 10 times larger.
+Similarly, the net effect of the disclosure of news, represented by and shown in Table 5, is 0.033 when the time window is 1 days, while the corresponding value for the direct effect is 0.600 (Table 4).
 Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms.
-Second, the difference between the post- and pre-news effects on suppliers is positive and significant only when the time window is 2 or 3 days, although the significantly positive difference between the post- and pre-news direct effects is observed in a longer run, up to the time window of 30 days.
-This implies that the net effect of the disclosure of news on suppliers lasts for only several days, a very short run, after the disclosure.
-Indirect effect on customers through global supply chains
-Similarly, we examine the effect on customer firms and show the results in Figure 8 and Table 5.
-In this analysis, although the pre- and post-news effects on customers show an inverted-U shape (panel [A] of Figure 8), as in the previous results, we do not find any positive and significant difference between the post- and pre-news effects (the upper part of Table 5).
-Therefore, in addition to the benchmark index of positiveness of each news article using the probabilities of positive and negative sentiment from the FinBERT (Figure 4), we employ an alternative positiveness index based on only the probability of positive sentiment shown by pink bars in Figure 3.
-As is clear from the comparison between Figures 3 and 4, the alternative index tends to evaluate the positiveness of news articles less than the benchmark index.
-Thus, using the alternative index, we can examine the effect of highly positive news.
-Then, we find a positive and significant difference between the post- and pre-news effects, or the net effect of the disclosure of highly positive news articles, on customers for the time windows of 2 and 3 days, as shown in the lower rows of Table 5.
-As in the case of the effect on suppliers, the net effect on customers is small and lasts only for several days.
-| 2 | 3 | 4 | 5 | 30 | 180 | 365 | |
-| Benchmark positiveness index using the positive and negative probabilities of news articles | |||||||
-| 0.0008 | -0.0007 | ||||||
-| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
-| Alternative positiveness index using only the positive probability of news articles | |||||||
-| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
+Second, the difference between the post- and pre-news effects on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, meaning that the effect of the disclosure of positive news about a firm on its suppliers shrinks by 88% in 5 days.
+By contrast, its effect on the firm’s own stock price shrinks from 0.600 to 0.157 by 74%.
+This comparison implies that the net effect of the disclosure of news on suppliers lasts shorter after the disclosure than the effect on own stock prices.
+Similarly, we examine the effect of positive and negative news articles on stock prices of client firms and illustrate the results in Panels (A) and (B) of Figure 7, respectively.
+Table 6 additionally presents the net effect of positive news, while the corresponding table for negative news can be found in the Supplementary Information (SI Table 2).
+These results are quite similar to those of the effect on suppliers shown in Figure 6 and Table 5.
+In summary, we find a positive (negative) and significant pre- and post-news effect of positive (negative) news and a positive (negative) and significant net effect of the disclosure of news.
+However, the effect of news about a firm on its clients is substantially smaller and lasts shorter than the effect on its own.
+| Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
+| value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.13 | 0.00 | 0.00 |
 Effect on Japanese firms
 We further use another sample that focuses on Japanese listed firms.
-The results for the direct effect of positive news articles on Japanese firms’ own stock prices shown in Figure 9 are qualitatively the same as and quantitatively similar to those using the sample of firms across the world shown in Figure 6.
-In addition, we examine the indirect effect of news articles about firms on stock prices of their suppliers and customers.
+The results for the direct effect of positive and negative news articles on Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5.
+One subtle difference is that in Figure 8 for Japan, the post-news effect of negative news using the time window of 2 days is larger in the absolute term than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite.
+This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
+In addition, we examine the indirect effect of news articles about Japanese firms on stock prices of their suppliers and clients.
 A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section).
 Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
-The results are presented in Figure 10 and Table 6.
-Panels (A) and (B) of Figure 10 for the effect on suppliers and customers, respectively, show several characteristics similar to each other but different from the previous results.
-First, although the pre- and post-news effect on stock prices of both suppliers and customers are mostly positive and significantly different from 0, as in the previous results, the pre-news effect is negative when the time window is 180 and 365 days, unlike the previous results.
-The negative pre-news effect in the long run implies that stock prices of firms positively mentioned in news do not necessarily grow faster long time before the news.
-In other words, the effect of news is less likely to be overestimated due to selection of potentially fast-growing firms.
-| Time window (days) | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
-| Effect on suppliers | |||||||
-| value | 0.21 | 0.14 | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 |
-| Effect on customers | |||||||
-| value | 0.00 | 0.00 | 0.44 | 0.00 | 0.00 | 0.00 | 0.00 |
-Second, the pre-news effect is positive and significant, implying diffusion of information before the disclosure of news that is consistent with the previous results.
-However, the difference between the post- and pre-news effects that can be interpreted as the net effect of the disclosure of news is negative in the short run, or in the time window of 2-5 days and turns to be positive in the long run (Table 6).
-This result is completely opposite to the previous ones using data for firms across countries (Tables 4 and 5).
-In other words, positive information about a firm diffuses through informal channels and raises the change rate of stock prices of the firm’s suppliers and customers before the information is disclosed in a news article.
-However, the disclosure of the information by a news article lowers the change rate of stock prices compared with that in the pre-news period, although the post-news effect is still positive.
-One possible reason for this is that investors who can receive information about firms through informal channels are more knowledgeable and recognize diffusion of performance through supply chains more accurately than other investors who rely on formal news.
-Therefore, once the knowledgeable investors receive positive information about a firm before the disclosure by news, they invest in the firm’s suppliers and customers.
-However, other investors react less to positive news about firms’ suppliers and customers.
-Finally, the pre- and post-news effects on suppliers and clients found in the sample of Japanese firms are larger than those in the sample of firms across the world.
-This finding, combined with the previous interpretation about knowledgeable investors, suggests that supply chain ties between suppliers and customers in Japan, known as keiretsu, are often stronger than in other countries, because they are often associated with shareholding relationships, information exchanges, technical assistance and research collaboration.[43, 44] Therefore, diffusion of performance through supply chains may be more prominent in Japan than in others, and this possibility is recognized by at least knowledgeable investors in Japan.
-Conclusion
-This study examines whether positive news about firms increases their stock prices and, moreover, whether it increases stock prices of the firms’ suppliers and customers, using a large sample of publicly listed firms across the world and another of Japanese listed firms.
-The level of positiveness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
-Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
-We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through informal channels.
-Positive news also raises stock prices of the firms’ suppliers and customers before its disclosure, confirming propagation of market values through supply chains.
-In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and customers than the pre-news effect.
-The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for informal information diffusion.
-However, the post-news effect on suppliers and customers in Japan is smaller than the pre-news effect, a result opposite to those from firms across the world.
+The results for the effect of positive and negative news on suppliers are demonstrated in Panels (A) and (B) of Figure 9, respectively, showing three notable differences from the previous results in Figure 6.
+First, the pre-news effect due to information diffusion through private channels is generally larger for Japanese firms than for others.
+For example, using the time window of 1 and 2 days, the pre-news effect of positive news on suppliers is 0.0090 and 0.0079 for Japan but 0.0029 and 0.0017 for others, respectively.
+Similarly, the post-news effect of negative news using the same windows is -0.0089 and -0.0066 for Japan and -0.0046 and -0.0025 for others, respectively.
+Second, the post-news effect of positive news is smaller than its pre-news effect (Panel [A] of Figure 9), although the post-news effect is always larger in size than the pre-news effect in the previous results.
+This finding implies that although positive information about a firm tends to diffuse actively through private channels, such as suggestions of financial companies, SNS, and word of mouth, before its disclosure by a news article and increase stock prices of suppliers of the firm.
+However, once the positive information is disclosed in a news article, the growth rate of the suppliers’ stock prices declines compared with that in the pre-news period.
+In other words, the net effect of the disclosure of positive news about a Japanese on stock prices of its suppliers is negative, although the previous results found the opposite.
+Finally, we do observe that the post-news effect is larger than the pre-news effect in the case of negative news, as in the previous results (Panel [B] of Figure 9).
+However, the pre-news effect of negative news for the time window of 1 day is smaller in the absolute term than that for the longer time window of 2-5 days, although the previous result for firms across the world showed the opposite (Panel [B] of Figure 6).
+This, combined with the large negative post-news effect for the time window of 1 day, implies that the disclosure of negative information about a firm by a news article affects stock prices of its suppliers more substantially than diffusion of the information through private channels.
+We further estimate the effect on clients and show the results in Figure 10.
+Compared with the previous results for firms across the world shown in Figure 7, we find that the three notable differences found in the effect on suppliers still hold in the effect on clients: (1) The average effect is generally larger for Japanese firms than for firms across the world; (2) The net effect of the disclosure of positive news about a firm on stock prices of its clients is negative; (3) The pre-news effect of negative news for the time window of 1 day is smaller than that for longer time windows.
+In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries.
+Specifically, we find that positive news propagates more before its disclosure than after in the Japanese data.
+Although we cannot provide evidence for any mechanism behind the differences, one possible reason is the difference in the quality of supply chain information.
+The supply chain data used for firms worldwide in this study relies on financial statements, whereas the data for Japan is based on a questionnaire-based survey.
+Although the former is automatically determined by criteria based on sales volume, the latter identifies important trade partners, which are not determined only by sales volume.
+Since investors who can receive information about firms through private channels are more knowledgeable about supply chains and recognize the diffusion of performance through supply chains more accurately than other investors who rely on news media, once the knowledgeable investors receive positive information about a firm before its disclosure by news media, they invest in the firm’s suppliers and clients.
+This results in a large increase in stock prices of the suppliers and clients before the disclosure of the news.
+However, other investors who rely on news media and do not understand the role of supply chains react less strongly to positive news after its disclosure.
+Another possible reason for this is the distinct nature of Japan’s supply chains, known as keiretsu.
+Links between suppliers and clients in these relationships are often stronger than those in other countries, as they are frequently associated with shareholding relationships, information exchanges, technical assistance, and research collaboration [67, 68].
+These strong relationships may explain why positive news about firms may have a greater impact on their suppliers and clients.
+Because investors recognize the strength of these ties and anticipate the transmission of performance through supply chains, news about firms tends to have a greater impact on the stock prices of their suppliers and clients in Japan compared to other countries.
+However, these two reasons cannot explain why negative news does not propagate before disclosure, which should be investigated in future research.
+5 Conclusion
+This study examines how positive and negative news about firms affects the change rate of their stock prices and, moreover, how it affects the change rate of stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms across the world and another of Japanese listed firms.
+The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
+Supply chains of firms across the world are identified mostly by financial statements and supplemented by other information on the website and news medias, while those of Japanese firms are taken from large-scale firm-level surveys.
+We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through private channels, such as SNS and word of mouth.
+Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains.
+In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect.
+The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels.
+However, the post-news effect on suppliers and clients in Japan can be smaller than the pre-news effect, a result opposite to those from firms across the world.
 This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
+The important opening question not addressed by this study is the significant heterogeneity of supply chain relationships, which can lead to varying impacts of news sentiment.
+For example, some suppliers are easily replaceable, while others are not; some trade relationships are long-term, potentially reflecting trust or sunk costs; additional links, such as shareholding or dual board membership, may exist; and the recognition of supply chain relationships by investors can vary.
+Examining the heterogeneity presents an important direction for future research.
+Data availability
+The data that support the findings of this study are available from FactSet Research Systems Inc.
+and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.
+but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available.
+Data are however available from the authors upon reasonable request and with permission of FactSet Research Systems Inc.
+and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.
 Acknowledgements
-The authors thank Yuito Nishi for his technical contribution.
+The authors thank Wataru Souma and Takayuki Yumoto for valuable discussions and Yuito Nishi for his technical contribution.
 Author contributions statement

### v4

@@
-Market Reaction to News Flows in Supply Chain Networks
+Stock Price Responses to Firm-Level News in Supply Chain Networks
 Abstract
-This study examines how positive and negative news about firms affects their stock prices and, moreover, how it affects stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms around the world and another of Japanese listed firms.
-The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
-Supply chains of firms across the world are identified mostly by financial statements, while those of Japanese firms are taken from large-scale firm-level surveys.
-We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through unofficial information channels.
-Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains.
-In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect.
-The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels.
-However, the post-news effect on suppliers and clients in Japan is smaller than the pre-news effect, which is the opposite result to non-domestic firms from around the world.
-keywords:
-Supply Chain, Natural Language Processing, Stock Market, Large -scale Data1 Introduction
+This study examines how positive and negative news about firms are associated with stock prices and whether these associations extend to suppliers and clients linked via supply chain relationships, using large samples of publicly listed firms worldwide and in Japan.
+News sentiment is measured using FinBERT, a natural language processing model fine-tuned for financial text, and supply chain links are identified from financial statements for global firms and from large-scale firm-level surveys for Japanese firms.
+We find that stock prices exhibit systematic associations with positive and negative news even before public disclosure.
+These associations are also observed for suppliers and clients before and after disclosure.
+In general, post-disclosure associations are larger than pre-disclosure associations, with the difference concentrated around the time of public news disclosure relative to the pre-disclosure period.
+However, for Japanese firms, the post-disclosure associations for suppliers and clients are smaller than the pre-disclosure associations, in contrast to the pattern observed for firms outside Japan.
+1 Introduction
 Firms interact with and influence each other through various types of networks.
-Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 2, 3] and knowledge networks through research collaboration.[4, 5, 6, 7, 8] Another type of network recently paid attention in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components.[9, 10] One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other.
-[11]
-Global supply chains have rapidly expanded because interconnected firms through supply chains can greatly benefit from the network.
-Most notably, supply chains can empower firms by enhancing efficiency of production processes and thus optimizing firms’ economic performance.[12, 13] For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity.[14] Moreover, it is often evidenced that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology.[15, 7, 16]
-Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability.[17, 18, 19] Such negative shocks examined in the literature include the Great East Japan Earthquake in 2011,[20] the massive flood in Thailand in 2011,[21] the COVID-19 pandemic from 2020 to 2022,[22], and the Russia-Ukraine war that started in 2022.[23]
-Therefore, positive or negative shocks propagate through supply chains, affecting performance of firms linked through supply chains.
-Because the corporate value of publicly listed firms is evaluated in the stock market,[24, 25, 26] a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms.
-For example, when Boeing 737 MAX crashed in 2019 and 2020, the stock price of Boeing, its supplier, such as GE and Allegheny Technologies, and its clients, such as American Airlines and Southwest Airlines, dropped sharply.[27, 28, 29, 30, 31].
-Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 negatively affected the stock price of Daihatsu, Toyota, and their suppliers and clients, such as Aisin.[32, 33]
+Major firm networks that are found to play a significant role in firms’ behaviors and performance in the literature include financial networks through ownership [1, 23, 27] and knowledge networks through research collaboration [30, 29, 10, 72, 34].
+Another type of network recently paid attention to in the academic literature and business and policy fields is supply chains formed through transactions of materials, parts, and components [15, 32].
+One reason for the growing attention is that supply chains have expanded globally, linking firms in the world within a small number of steps with each other [6].
+Global supply chains have rapidly expanded because firms interconnected through supply chains can greatly benefit from the network.
+Most notably, supply chains can empower firms by enhancing the efficiency of production processes and thus optimizing firms’ economic performance [11, 5].
+For instance, by procuring and manufacturing different goods in various locations based on each location’s comparative advantages and cost-effectiveness, firms can streamline their operations as a network, ultimately leading to reductions in production costs and improvements in productivity [18].
+Moreover, it is often evident that productivity spills over through supply chains from upstream suppliers to downstream client firms because of high-quality materials, parts, components, and services and from downstream to upstream firms because of learning of knowledge and technology [55, 72, 38].
+Conversely, supply chains can be a channel of negative shocks that aggravate firm performance, particularly once supply chains are disrupted due to unexpected events, such as natural disasters and geopolitical instability [9, 35, 7].
+Such negative shocks as examined in the literature include the Great East Japan Earthquake in 2011 [49], the massive flood in Thailand in 2011 [31], the COVID-19 pandemic from 2020 to 2022 [17], and the Russia-Ukraine war that started in 2022 [54].
+Accordingly, prior studies show that positive or negative shocks propagate through supply chains, affecting the performance of firms linked through supply chains [7].
+Because the corporate value of publicly listed firms is evaluated in the stock market [52, 71, 12], a shock to a firm is expected to affect not only its own stock price but also the stock prices of its suppliers and client firms.
+For example, following the Boeing 737 MAX crashes in 2019 and 2020, stock prices of Boeing, some of its suppliers such as GE and Allegheny Technologies, and some of its clients such as American Airlines and Southwest Airlines declined sharply [13, 14, 22, 48, 62].
+Also, a misconduct by Daihatsu Motor Co., a Japanese automobile manufacturer affiliated with Toyota, in 2023 was accompanied by stock price declines for Daihatsu, Toyota, and related firms including suppliers and clients such as Aisin [64, 63].
 These examples show that the shock can propagate through supply chains and affect stock prices of firms linked through supply chains.
-Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices[34, 35, 36], there are not so many studies that investigated the diffusion of shocks to other firms’ stock prices through supply chains.
+Although relatively a lot of previous studies have examined the direct impact of shocks to firms on their own stock prices [44, 69, 68], there are not so many studies that have investigated the diffusion of shocks to other firms’ stock prices through supply chains.
+Cohen and Frazzini [16] show that information about a firm’s customers predicts future stock returns of its suppliers, providing early evidence that stock price reactions propagate through supply chain relationships.
+More recently, Le Tran and Coqueret [42] show that firm-level news shocks, particularly ESG-related news, affect not only the focal firm’s stock returns but also the returns of its suppliers and clients, providing direct evidence of news-driven spillovers along supply chain relationships.
+Complementing this evidence, Agarwal et al.
+[2] demonstrate that investor attention, measured through online co-search behavior, facilitates information diffusion across supply chain partners and contributes to predictable stock return comovements.
 Kim and Wagner [40] analyzed 315 cases reported by the Wall Street Journal regarding supply chain risks, which also included mentions of 69 suppliers and 246 clients.
 Pandit et al.
-utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyzed their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT[38].
+utilized a larger dataset of quarterly earnings announcements, which differ from the news reports considered in our study, and analyze their propagation across 1,518 suppliers and 4,010 clients in COMPUSTAT [57].
 In addition, similar studies focusing on China are discussed [45, 46].
 However, to the best of our knowledge, no study has examined large-scale data spanning more than a decade and covering tens of thousands of firms.
-To fill this gap, this study examines this possible diffusion, using two large firm-level samples, one for firms across the world and the other for Japanese firms.
+To fill this gap, this study examines whether news sentiment is associated with stock price changes of firms and of their suppliers and clients in the pre- and post-disclosure periods, using two large firm-level samples, one for firms across the world and the other for Japanese firms.
 Specifically, we test whether the level of positive and negative sentiment about a firm in a news article affects the change rate of the stock price of the firm and its suppliers and clients before and after the disclosure of the news.
-We contribute to the literature by utilizing large-scale data in many respects.
-First, our firm-level data covers most publicly listed firms in the world and in Japan and identify detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys.
-Second, our data on news articles are taken from the NewsArchive of Thomson Reuters and cover more than 20 million articles.
-Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive news.
-Although we have mainly focused on how the shock affects stock prices through supply chains, it is evident that this study also relates to business applications and researches that predict changes of stock prices based on the news.
+Beyond supply-chain-specific mechanisms, a growing literature examines how the content of news itself is quantified and linked to stock market reactions.
+Early studies quantified news content using dictionary-based and word-frequency approaches to capture investor sentiment and market reactions.
+For example, Tetlock [70] construct a sentiment measure based on word categories in financial news and show that media pessimism predicts short-run market returns and trading volume.
+Relatedly, Mizuno et al.
+[53] quantify the novelty and topicality of business news using word-level similarity measures and demonstrate that such text-based characteristics are strongly associated with subsequent stock market activity.
 In this field, neural networks are predominantly used.
-For example, hedge funds, such as Bridgewater Associates utilizes the neural networks for predictions[41].
+For example, hedge funds, such as Bridgewater Associates, utilize neural networks for predictions [8].
 In academia, the transformer [73] is widely used [67, 25].
 In addition to the transformer, there is a long history of the usage of neural networks to predict stock prices, such as long short-term memory [37, 41, 20, 21, 39, 43, 65, 26], convolutional neural networks [33], or other deep learning methods [47, 58, 66, 56].
-We use sentiment analysis to determine the positiveness and negativeness of each news article[58, 59] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT)[60] has garnered the most acclaim[35].
+We use sentiment analysis to determine the positiveness and negativeness of each news article [51, 28] and for this purpose, specifically, Bidirectional Encoder Representations from Transformers (BERT) [19] have garnered the most acclaim [69].
 We take advantage of FinBERT [4], a fine-tuned model of BERT particularly for financial information.
+To advance this line of research from market- or firm-level settings to supply-chain networks and cross-country comparisons, it is essential to combine large-scale news data with comprehensive firm-level linkages and stock price information.
+We contribute to the literature by utilizing large-scale data in many respects.
+First, our firm-level data covers most publicly listed firms in the world and in Japan and identifies detailed supply chains among them using financial statements, news articles, websites, and firm-level surveys.
+Second, our data on news articles are taken from the News Archive of Thomson Reuters and cover more than 20 million articles.
+Third, we employ daily stock prices of all listed firms to examine both short- and long-run effects of positive and negative news.
 2 Data
-This study utilizes four major data sources.
-First, our information on news relies on news articles written in English included in the NewsArchive provided by Thomson Reuters for the period 2003-2016.
-The information in each news article includes the time, headline, main text, Refinitiv Identification Code (a ticker-like code) of firms mentioned in the article, language, etc.
+2.1 News article
+Our information on news relies on news articles written in English included in the News Archive provided by Thomson Reuters for the period 2003–2016.
+In the News Archive, multiple news articles may be filed for a single event.
+Typically, the first article is an alert, which is a short message containing only essential facts.
+This is followed by a newsbreak, usually released within 5–20 minutes after the alert.
+A newsbreak consists of a headline, which is similar to that of the alert, and two to four paragraphs of body text.
+Subsequent updates may follow the newsbreak after an additional 20–30 minutes, providing further information.
+In this study, we use the newsbreak associated with each event.
+Although we are interested in the timing of public disclosure, alerts often do not include sufficient details about the related firms, whereas newsbreaks do.
+The time difference between alerts and newsbreaks is typically less than 20 minutes, which is negligible given that our analysis is conducted at the daily frequency.
+Accordingly, the release time of the newsbreak is aligned with the closest market closing time of the focal stock market, and the corresponding trading day is defined as day zero in the subsequent analyses.
+The newsbreak includes the time, headline, main text, Refinitiv Identification Codes (RICs) of one or more listed firms mentioned in the article, the language, and other metadata.
 As we will explain in detail later, our sentiment analysis, which determines the degree of positivity or negativity of each news article about the firms mentioned, utilizes the main text.
-The average number of words in an article is 243, whereas its median is 136 and maximum is 3,911.
+The average number of words in an article is 243, whereas its median is 136 and its maximum is 3,911.
 Although the total number of news articles in the News Archive for the period examined is 20,803,561, we focus on 3,447,425 that mention any firm publicly listed in 105 stock markets in the world.
-Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33.
-(Figure 1).
+Among these news articles, 90% mention three or fewer firms, while the maximum number of firms mentioned in a single article is 33 (Figure 1).
 The number of news articles that mention a particular firm during the period 2003-2016 is quite skewed.
 55.6% of firms were never mentioned, while 0.1% were mentioned more than 10,000 times (Figure 2).
-Second, we utilize daily stock prices of currently listed firms in stock markets in the world taken from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, using its Application Programming Interface (API).
-We drop firms listed in the past but not currently from our sample, because their stock prices are not available using the API.
-We particularly define the daily stock price of each firm as its closing price.
+Stock price
+We utilize daily stock prices of currently listed firms in global stock markets, obtained from the Eikon database provided by Refinitiv, one of the largest providers of financial markets data and infrastructure, through its Application Programming Interface (API).
+Firms that were listed in the past but are no longer listed are excluded from the sample, as their stock price data are not available through the API.
+We define the daily stock price of each firm as its closing price.
+Accordingly, the stock price data include Refinitiv Identification Codes (RICs), dates, and closing prices.
 In addition to stock prices of each firm, we utilize the Refinitiv index for each stock market that indicates the average of stock prices in the market to control for the market-specific time trend.
 The Refinitiv index is available for 105 major markets that cover 92% of listed firms worldwide.
-Third, to identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023.
-The FactSet data include major clients of each firm reported in its financial statement.
+Supply chain
+To identify global supply chains among publicly listed firms in the world, we rely on data from FactSet that cover a number of firms across the world for the period 2003–2016 obtained as of January 18, 2023.
+The FactSet data include major clients of each firm reported in its financial statements.
 The US, Japanese, and many other governments require every listed firm to disclose its major clients with which the sales are more than 10% of its total sales, at least in the period examined in this study [24].
@@
 The number of firms covered in the FactSet data increased rapidly over time, because the dataset has been constructed recently and thus ignores firms in the distant past.
-In a later section, the recorded supply chain links at the time are considered to estimate the impact of the news on supply chains.
+In a later section, the recorded supply chain links at the time are used to examine news-related stock price changes along supply chain relationships.
+The FactSet supply chain database does not cover the universe of firms worldwide.
+Instead, it identifies major customer–supplier relationships disclosed in financial statements and related sources, which are typically limited to economically significant partners.
+As a result, its coverage is intentionally limited, especially in earlier years.
+FactSet uses its own identifier system, referred to as the FactSet Security Identifier.
+We convert these identifiers to Refinitiv Identification Codes (RICs) in order to link the FactSet data with news and stock price data.
 | 2003 | 2004 | 2005 | 2006 | 2007 | 2008 | 2009 | |
@@
 | Maximum outdegree | 40 | 54 | 79 | 79 | 75 | 109 | 140 |
-Finally, to identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan.
-Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms of each firm.
+To identify domestic supply chains within Japan, we employ data collected by Tokyo Shoko Research (TSR) for the period 2008–2016, particularly its Company Linkage Database that covers supply chains of more than one million firms in Japan.
+Supply chains are captured by annual firm-level surveys by TSR that request each firm to report up to 23 suppliers and client firms for each surveyed firm.
 Apparently, many firms are linked with more than 23 suppliers and clients, and thus their links cannot be fully identified by their own responses.
@@
 Since those firms are relatively large, their partners report them as important suppliers or clients.
+Therefore, although the numbers of suppliers and clients are limited to 23 in the reports, large firms can have many partners.
+Accordingly, the maximum number of suppliers and clients is 12,729 and 12,016, respectively, in 2016, providing broad coverage of supply chains of Japanese firms.
+The TSR database provides exceptionally broad coverage of firm-level relationships, including both listed and unlisted firms.
+In this study, however, we restrict our analysis to relationships involving listed firms and links between such firms, in order to ensure consistency with market-based information.
+Listed firms in the TSR database are identified by corporate IDs and converted into Refinitiv Identification Codes (RICs).
 The number of firms and their supply-chain links and the corresponding number for listed firms from 2008 to 2016 are presented in Table 2.
@@
 (2018) [19].
-Specifically, FinBERT uses Reuters’ TRC2-financial that consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank that consists of 4,845 English sentences taken from financial news in the LexisNexis database.
-From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three about the firms mentioned in the article.
+Specifically, FinBERT uses Reuters’ TRC2-financial, which consists of 1.8 million news articles between 2008 and 2010 and fine-tunes BERT for financial sentiment classification using Financial Phrasebank, which consists of 4,845 English sentences taken from financial news in the LexisNexis database.
+From each news article, FinBERT generates softmax outputs for three labels, i.e., positive, neutral, and negative, which indicate the weights of the three sentiments about the firms mentioned in the article.
 Examples are shown in Table 3.
@@
 Both panels indicate that the probability of neutral sentiment is high for many news articles followed by that of negative sentiment while news articles with a large probability of positive sentiment are fewer.
+From the three sentiment probabilities generated by FinBERT (positive, neutral, and negative), we construct sentiment measures by excluding the neutral component and normalizing the remaining positive and negative probabilities.
+Specifically, the positive sentiment index is defined as
+and the negative sentiment index as
+where and denote the probabilities of positive and negative sentiment, respectively, for news article mentioning firm on day .
+This normalization focuses the analysis on the relative balance between positive and negative information and avoids ambiguity associated with neutral sentiment.
 | Sentiment | |||
@@
 3.1 Regression analysis
-Direct effect
-To estimate the effect of positive and negative sentiment in news articles about particular firms on their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis.
-Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window are affected by the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure.
+Own-firm associations
+To examine the association between positive and negative sentiment in news articles about particular firms and their own stock prices, we apply the probabilities of positive and negative sentiment constructed above to regression analysis.
+Specifically, we focus on firms mentioned by any news article and examine whether the change rate of their stock prices in a time window is associated with the level of positiveness or negativeness of the article after the article is disclosed, compared with its change rate before the disclosure.
 Suppose that firm listed in stock market is mentioned by news article on day .
-Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article disclosed on day in time window , or and , respectively, by
+Setting the time window at , such as 1, 5, and 30 days, we define the average daily percentage change of the stock price of firm before and after news article was disclosed on day in time window , or and , respectively, by
 where is the average stock price of firm between day and .
@@
 where T is either pre or post, and and are indicator variables that take a value of one if represents the percentage change in the stock price of firm before or after news article is disclosed on day , respectively.
-represents the level of either positive or negative sentiment in news article generated by FinBERT (Data section).
+represents the level of positive or negative sentiment in news article generated by FinBERT and corresponds to the sentiment indices and defined in the Methods section.
 is the percentage change of the Refinitiv index for each stock market that controls for the overall time trend of stock prices in the market, and is sector dummies.
 Sectors are defined by the North American Industry Classification System at the sub-sector level for the sample of firms across the world and by the Japan Standard Industrial Classification at the two-digit level for the sample of Japanese firms.
-Accordingly, indicates the average effect of the positiveness or negativeness of a news article on the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors.
-By contrast, indicates any prior “effect” of the news article on the change rate of the stock price, including the possible effect of the information in the news article that diffuses in the market not through the news but through unofficial information channels before the disclosure of the news.
-Such unofficial information channels include suggestions by financial companies to their clients, posts in Social Networking Services (SNS) by influencers, and word of mouth among investors.
+Accordingly, indicates the average association between the positiveness or negativeness of a news article and the average change rate of the stock price of the firm mentioned in the article after it is disclosed in the time window of days, controlling for the overall trend in the market and sector-specific unobservable factors.
+By contrast, captures any prior association between the news article and the change rate of the stock price before disclosure.
+This term may reflect a combination of factors that are not separately identified in this analysis, including selection into news coverage and pre-existing price dynamics.
 In addition, may include the endogeneity bias reflecting the possibility that the growth rate of the stock price of a firm that is positively mentioned by any news article is intrinsically higher than that of a firm negatively mentioned.
-Therefore, the net effect of the disclosure of the information by the news article on the change rate of the stock price in time window should be given by [66].
-We particularly estimate equation (3) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run effect of positive news articles on stock prices.
-We hypothesize that a positive (negative) news article would accelerate (deteriorate) the change rate of the stock price of the firm mentioned in the article in the short-run, while the effect would become smaller and converge to 0 in the longer run.
+Accordingly, the difference between the post- and pre-news coefficients, , summarizes changes in stock price growth around the time of news disclosure [36], relative to the pre-disclosure period.
+This difference compares post-disclosure price movements with those observed before disclosure, where pre-news movements may reflect the incorporation of information prior to public disclosure.
+We particularly estimate equation (5) for 1, 2, 3, 4, 5, 30, 180, and 365 to examine both the short- and long-run associations between positive news articles and stock prices.
+We hypothesize that a positive (negative) news article is associated with higher (lower) stock price changes in the short-run, with weaker associations in the longer run.
 It should be emphasized that this analysis is comparing firms positively mentioned in news articles with those negatively mentioned, rather than comparing firms positively mentioned and those not mentioned by news articles.
 We focus on the sub-sample of firms mentioned by news articles and ignore those not mentioned, because the two types of firms are most likely to be intrinsically different from each other to a great extent so that comparing the two may be associated with large endogeneity biases.
-Indirect effect through supply chains
-We further examine the effect of positive or negative sentiment about firms in news articles on their suppliers and clients, using similar frameworks.
-Specifically, when we focus on the upstream effect of news articles about firms on their suppliers, in equation (3) is replaced with the probability of positive or negative sentiment in the news article that mentions not firm but any of ’s client firms.
-Alternatively, when we examine the downstream effect of clients’ news on their clients, is replaced with the positive or negative index of any of the suppliers of firm .
+Associations along supply chains
+We further examine the association between positive or negative sentiment about firms in news articles and stock price changes of their suppliers and clients, using similar frameworks.
+Specifically, when we focus on upstream associations between news articles about firms and their suppliers, the variable in equation (5) is replaced with the positive or negative sentiment indices from news articles that mention any of firm ’s client firms.
+Alternatively, when we examine the downstream associations between news articles about firms and their clients, is replaced with the positive or negative indices of any of the suppliers of firm .
 When a firm mentioned by a news article has more than one supplier or client, as shown in Tables 1 and 2, we include all the suppliers and clients in our sample.
-In this analysis, we compare stock prices of suppliers or clients of firms mentioned positively in news articles with those of suppliers or clients of firms mentioned negatively, as in our analysis on the direct effect on own stock prices.
-We hypothesize that the effect of news articles on stock prices diffuses through supply chains.
-Precisely, a positive news article about a client firm would accelerate the growth in the stock price of its suppliers, in addition to its own stock price, because the market predicts an increase in demand for products of the suppliers associated with higher performance in the client firm.
-Also, positive news about a supplier would promote growth of its clients because of possible productivity growth of the supplier.
-As in the case of the direct effect, the indirect effect on suppliers and clients would be smaller in the long run than in the short run.
-4 Results and Discussion
-Direct effect of news articles on listed firms across the world
-We start with estimating how the positiveness of news articles about firms affect percentage changes in their stock prices before and after the disclosure of news, applying equation (3) to the sample of publicly listed firms across the world mentioned by any news article.
+We hypothesize that news sentiment about a firm is associated with stock price changes of its suppliers and clients.
+More precisely, news sentiment about a client firm is associated with stock price changes of its suppliers, in addition to those of the client firm itself, while news sentiment about a supplier firm is associated with stock price changes of its clients.
+These associations may reflect, for example, investors’ expectations about changes in demand faced by suppliers or about the performance of downstream firms, although such mechanisms are not directly tested in this study.
+As in the case of own-firm associations, these associations are expected to be stronger over shorter horizons and weaker over longer horizons.
+4 Results
+Own-firm stock price associations for listed firms across the world
+We start by examining how the positiveness of news articles about firms is associated with percentage changes in their stock prices before and after the disclosure of news, applying equation (5) to the sample of publicly listed firms across the world mentioned by any news article.
 Results are presented in Table 4 and Figure 5.
+Then, we show the results for the negativeness.
 | Dependent Variable: | Daily percentage change of the stock price of the firm mentioned in each news article in the pre- or post-news period | |||||||
@@
 | Number of observations | 9,409,978 | 9,543,586 | 9,562,439 | 9,565,054 | 9,562,897 | 9,460,176 | 8,986,209 | 8,546,395 |
-In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that the probability of positive sentiment in a news article about a firm is positively correlated with the percentage change in the firm’s stock price from 2 days to 1 day before the disclosure of the news.
+In column (1) of Table 4, the estimated is 0.323 and highly significant, indicating that a higher value of the positive sentiment index in a news article about a firm is associated with the percentage change in the firm’s stock price from 2 days to 1 day before the release of the news.
 Since the positiveness measure, in equation (5), ranges from 0 to 1 (Figure 3), the value of the coefficient means that the change rate of the stock price of a firm mentioned in a news article most positively () is 0.323% points higher than that of a firm mentioned with no positive sentiment () before the disclosure of the news.
 There are two possibilities for this positive coefficient.
-First, positive information included in news articles diffuses to stock markets partially through private channels before the formal disclosure of news, generating the pre-news effect.
-Second, stock prices of firms positively mentioned by news articles grow faster intrinsically than those of firms with no positive mention, leading to endogeneity biases.
-Column (1) of Table 4 further finds that the estimated is 0.923, indicating that the stock price of a firm mentioned in a news article most positively grows faster by 0.923% points than that of a firm mentioned with no positive sentiment after the news disclosure.
+First, positive information included in news articles may be incorporated into stock prices prior to formal disclosure through non-public channels, contributing to pre-news price movements.
+Second, stock prices of firms positively mentioned by news articles may exhibit higher intrinsic growth than those of firms with no positive mention, leading to endogeneity biases.
+In the following analysis, we do not attempt to disentangle these two channels.
+Instead, we focus on the difference between post- and pre-news coefficient estimates, which summarizes the additional change in stock prices around public news disclosure relative to pre-existing price movements.
+Column (1) of Table 4 further shows that the estimated is 0.923, indicating that, in the post-news period, firms mentioned with positive news sentiment experience stock price growth that is higher by 0.923 percentage points than firms mentioned with no positive sentiment.
 The second and third rows from the bottom show and the value from a test for the null hypothesis that the difference is 0.
-In column (1), the value is close to 0, and thus we reject the null hypothesis.
-This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, positive news raises stock prices of former firms more by 0.6% points after its disclosure.
-These results in column (1) for the time window of 1 day generally hold in columns (2)-(5) for time windows of 2-5 and 30 days, while the post- and pre-news effects and the difference between the two decrease as the time window becomes wider.
-By contrast, the difference between the post- and pre-news effects is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size.
+In column (1), the value is close to 0, indicating that the difference is statistically different from zero.
+This result suggests that although stock prices of firms mentioned by news positively grow faster than those of firms mentioned with no positive sentiment even before the disclosure of news, stock prices of firms with positive news sentiment are higher by an additional 0.6% after its disclosure.
+These results in column (1) for the time window of 1 day generally hold in columns (2)-(6) for time windows of 2-5 and 30 days, while the post- and pre-news coefficient estimates and the difference between the two decrease as the time window becomes wider.
+By contrast, the difference between the post- and pre-news coefficient estimates is negative when the time window is 180 and 365 days, or in the long run, although the difference is quite small in size.
 This result implies that the intrinsic difference between firms mentioned positively and those with no positive mention in news articles is negligible.
 Panel (A) of Figure 5 graphically demonstrates the argument above.
-The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news effect for different time windows, whereas the right half shows the post-news effect.
+The left half of the figure where the value on the horizontal axis is negative shows the point estimate of the pre-news coefficient for different time windows, whereas the right half shows the post-news coefficient.
 Although the confidence intervals are added in the figure, they are too small to be recognized.
 The estimated coefficients show an inverted-U shape where the coefficients are close to 0 on the left and right edges.
-This shape suggests the followings.
-The change rate of the stock price of firms positively mentioned by news articles is not intrinsically different from that of firms negatively mentioned long before the disclosure of the news.
+This pattern indicates the following.
+The change rate of the stock price of firms mentioned with positive news sentiment is not intrinsically different from that of firms mentioned with no positive news sentiment long before the disclosure of the news.
 However, the change rate of the stock price of positively mentioned firms began to rise compared to firms with no positive mention 30 days before the disclosure, peaking one day after the disclosure.
 Then, the difference in the change rates of stock prices between positively mentioned firms and those with no positive mention gradually decreases over time, converging to nearly zero 180 days after the disclosure.
-The effect of the probability of negative sentiment in each news article shown in Panel (B) of Figure 5 is consistent with the effect of the probability of positive sentiment in Panel (A).
-The effect of negative sentiment in a news article is negligible 180 days or more before its disclosure but becomes negative and significant 30 days or fewer before the disclosure, implying information diffusion through private channels.
-After the disclosure, the negative effect becomes larger in the absolute term, suggesting a negative net effect of the disclosure of negative news on stock prices, in addition to the negative effect through private channels.
-Indirect effect on suppliers and clients through global supply chains
-Next, we examine whether positive news about a firm affects stock prices of the firm’s suppliers and clients through global supply chains by estimating equation (3) where the positiveness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
-Panels (A) and (B) of Figure 6 are the graphical presentation of the results for the effect of positive and negative news about firms, respectively, on their suppliers.
-Table 5 particularly shows the difference between the post- and pre-news effects of positive news and its values.
-The corresponding table for the effect of negative news is omitted here but is available in the Supplementary Information (SI Table 1).
-The results shown in the figure and table are similar to the results for the direct effect on own firms’ stock prices (Figure 5 and Table 4).
-The post- and pre-news effects of positive news are positive and significant, and their difference is positive and significant when the time window is narrow.
-These results suggest that positive news about firms raises stock prices of the firms’ suppliers before the disclosure of the news due to information diffusion through private channels and raises them more after the disclosure due to the additional effect of news medias.
-The effect of negative news articles is opposite in the direction and similar in size to the effect of positive ones, indicating the same conclusion for negative news.
+The pattern of coefficients for negative sentiment shown in Panel (B) of Figure 5 is qualitatively consistent with the pattern of positive sentiment shown in Panel (A).
+The corresponding detailed regression results are reported in the Supplementary Information (SI Table 3).
+The coefficients are negligible 180 days or more before its disclosure but become negative and significant 30 days or fewer before the disclosure.
+After disclosure, the magnitude of the coefficients increase further.
+Associations for suppliers and clients along global supply chains
+Next, we examine whether news sentiment about a firm is associated with stock prices of the firm’s suppliers and clients along global supply chains by estimating equation (5) where the positiveness or negativeness index of news article about firm on day , , is replaced with the index for any supplier or client of firm .
+Panels (A) and (B) of Figure 6 present coefficient estimates for positive and negative news sentiment about firms and the associated stock price changes of their suppliers.
+Table 5 reports the difference between post- and pre-news coefficient estimates for positive sentiment, with corresponding results for negative sentiment reported in SI Table 1.
+Detailed regression results are reported in SI Tables 4 and 5.
+For suppliers, both pre- and post-news coefficient estimates are positive and statistically significant for positive news sentiment, with a larger post–pre difference for shorter time windows.
+The corresponding coefficient estimates for negative sentiment exhibit the opposite sign and similar magnitudes.
+Overall, the magnitude of supplier-level coefficient estimates is substantially smaller than that observed for own-firm stock prices.
 | Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
 | value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.27 | 0.00 | 0.00 |
-However, there are two notable differences between the direct effect on own firms and the indirect effect on suppliers.
-First, the post- and pre-news effect on suppliers and their difference is substantially smaller than the direct effects and their difference.
-For example, the pre-news effect on suppliers for the time window of 1 days is 0.0293 (Figure 6), while the corresponding direct effect is 0.323 (Figure 5), more than 10 times larger.
-Similarly, the net effect of the disclosure of news, represented by and shown in Table 5, is 0.033 when the time window is 1 days, while the corresponding value for the direct effect is 0.600 (Table 4).
-Therefore, we conclude that although positive news articles about firms affect stock prices of their suppliers, the effect on suppliers is less than 10% of the direct effect on own firms.
-Second, the difference between the post- and pre-news effects on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, meaning that the effect of the disclosure of positive news about a firm on its suppliers shrinks by 88% in 5 days.
-By contrast, its effect on the firm’s own stock price shrinks from 0.600 to 0.157 by 74%.
-This comparison implies that the net effect of the disclosure of news on suppliers lasts shorter after the disclosure than the effect on own stock prices.
-Similarly, we examine the effect of positive and negative news articles on stock prices of client firms and illustrate the results in Panels (A) and (B) of Figure 7, respectively.
-Table 6 additionally presents the net effect of positive news, while the corresponding table for negative news can be found in the Supplementary Information (SI Table 2).
-These results are quite similar to those of the effect on suppliers shown in Figure 6 and Table 5.
-In summary, we find a positive (negative) and significant pre- and post-news effect of positive (negative) news and a positive (negative) and significant net effect of the disclosure of news.
-However, the effect of news about a firm on its clients is substantially smaller and lasts shorter than the effect on its own.
+However, there are two notable differences between the own-firm coefficients and those for suppliers.
+First, the post- and pre-news coefficients for suppliers and their difference are substantially smaller than the corresponding own-firm coefficients and their difference.
+For example, the pre-news coefficient for suppliers for the time window of 1 day is 0.0293 (SI Table 4), while the corresponding coefficient for own firms is 0.323 (Table 4), more than 10 times larger.
+Similarly, the difference between the post- and pre-news coefficients, , is 0.033 for suppliers when the time window is one day (Table 5), while the corresponding difference for own firms is 0.600 (Table 4).
+Quantitatively, although positive news articles about firms are associated with stock prices of their suppliers, the corresponding coefficients for suppliers are less than 10% of those for own firms.
+Second, the difference between the post- and pre-news coefficients on suppliers is 0.033 when the time window is 1 day but becomes 0.004 when it is 5 days, indicating that the post-pre difference in coefficients for suppliers declines by 88% in 5 days.
+By contrast, the coefficients for own firm decline from 0.600 to 0.157 by 74%.
+This comparison shows that the difference between the post- and pre-news coefficients for suppliers diminishes more quickly after disclosure than the corresponding difference for firms’ own stock prices.
+As in the analysis for firms’ own stock prices, negative sentiment shows a pattern similar to positive sentiment for suppliers (SI Table 5).
+Both the post- and pre-news coefficients for suppliers, and their difference, are substantially smaller than the corresponding coefficients for firms’ own stock prices, and the post–pre difference for suppliers diminishes more quickly after disclosure.
+Similarly, we examine coefficient estimates for positive and negative news sentiment in relation to stock prices of client firms and present the results in Panels (A) and (B) of Figure 7, respectively.
+We also show the difference between the post- and pre-news coefficient estimates for positive (Table 6) and negative (SI Table 2).
+Detailed regression results for clients are reported in SI Table 6 for positive sentiment and in SI Table 7 for negative sentiment.
+These results exhibit the same qualitative pattern as those for suppliers shown in Figure 6 and Table 5.
+Specifically, we find positive (negative) and significant pre- and post-news coefficients for positive (negative) news sentiment, as well as a positive (negative) and significant post–pre difference.
+However, the coefficients for clients are substantially smaller and decline more rapidly than those for own firms.
 | Time window (days) | 1 | 2 | 3 | 4 | 5 | 30 | 180 | 365 |
 | value | 0.00 | 0.00 | 0.00 | 0.00 | 0.00 | 0.13 | 0.00 | 0.00 |
-Effect on Japanese firms
+Japanese firms
 We further use another sample that focuses on Japanese listed firms.
-The results for the direct effect of positive and negative news articles on Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5.
-One subtle difference is that in Figure 8 for Japan, the post-news effect of negative news using the time window of 2 days is larger in the absolute term than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite.
+The results for the own-firm coefficients of positive and negative news sentiment for Japanese firms’ own stock prices shown in Figure 8 are qualitatively and quantitatively similar to those using the sample of firms across the world shown in Figure 5.
+The detailed regression results are shown in SI Tables 8 and 9.
+One subtle difference is that in Figure 8 for Japan, the post-news coefficient for negative news sentiment using the time window of 2 days is larger in absolute terms than that using the time window of 1 day, while Figure 5 for firms in the world shows the opposite.
 This finding suggests that the market reaction to negative news in Japan is slower than that to positive news in Japan and to positive and negative news in the world.
-In addition, we examine the indirect effect of news articles about Japanese firms on stock prices of their suppliers and clients.
-A notable difference between this analysis on Japanese firms and the previous one on firms across the world is that more detailed domestic supply chains are identified based on firm-level surveys in the former than in the latter that mostly relies on financial statements (Data Section).
-Therefore, the average number of supply chain links for each Japanese firm in 2016 is 6.4, while that for each firm in the world is 3.3 (Tables 1 and 2).
-The results for the effect of positive and negative news on suppliers are demonstrated in Panels (A) and (B) of Figure 9, respectively, showing three notable differences from the previous results in Figure 6.
-First, the pre-news effect due to information diffusion through private channels is generally larger for Japanese firms than for others.
-For example, using the time window of 1 and 2 days, the pre-news effect of positive news on suppliers is 0.0090 and 0.0079 for Japan but 0.0029 and 0.0017 for others, respectively.
-Similarly, the post-news effect of negative news using the same windows is -0.0089 and -0.0066 for Japan and -0.0046 and -0.0025 for others, respectively.
-Second, the post-news effect of positive news is smaller than its pre-news effect (Panel [A] of Figure 9), although the post-news effect is always larger in size than the pre-news effect in the previous results.
-This finding implies that although positive information about a firm tends to diffuse actively through private channels, such as suggestions of financial companies, SNS, and word of mouth, before its disclosure by a news article and increase stock prices of suppliers of the firm.
-However, once the positive information is disclosed in a news article, the growth rate of the suppliers’ stock prices declines compared with that in the pre-news period.
-In other words, the net effect of the disclosure of positive news about a Japanese on stock prices of its suppliers is negative, although the previous results found the opposite.
-Finally, we do observe that the post-news effect is larger than the pre-news effect in the case of negative news, as in the previous results (Panel [B] of Figure 9).
-However, the pre-news effect of negative news for the time window of 1 day is smaller in the absolute term than that for the longer time window of 2-5 days, although the previous result for firms across the world showed the opposite (Panel [B] of Figure 6).
-This, combined with the large negative post-news effect for the time window of 1 day, implies that the disclosure of negative information about a firm by a news article affects stock prices of its suppliers more substantially than diffusion of the information through private channels.
-We further estimate the effect on clients and show the results in Figure 10.
-Compared with the previous results for firms across the world shown in Figure 7, we find that the three notable differences found in the effect on suppliers still hold in the effect on clients: (1) The average effect is generally larger for Japanese firms than for firms across the world; (2) The net effect of the disclosure of positive news about a firm on stock prices of its clients is negative; (3) The pre-news effect of negative news for the time window of 1 day is smaller than that for longer time windows.
+In addition, we examine coefficients for news sentiment about Japanese firms on stock prices of their suppliers and clients.
+Detailed regression results for suppliers are reported in SI Tables 10 and 11, and those for clients are reported in SI Tables 12 and 13.
+A notable difference between the Japanese supply chain data and the global supply chain data is that the former provides dense coverage of domestic supply chain links based on firm-level surveys, whereas the latter identifies major customer–supplier links primarily from financial statements (see Data Section).
+Consistent with this difference in coverage, the average number of supply chain links per listed firm in 2016 is 6.4 in Japan and 3.3 in the global sample (Tables 1 and 2).
+The coefficient estimates for positive and negative news sentiment in relation to suppliers’ stock prices are shown in Panels (A) and (B) of Figure 9, respectively, revealing three notable differences from the corresponding global results shown in Figure 6.
+First, the pre-news coefficients, which are larger in magnitude prior to disclosure, are generally larger for Japanese firms than for firms in the global sample.
+For example, for the one- and two-day time windows, the pre-news coefficients for positive news on suppliers are 0.0090 and 0.0079 in Japan, compared with 0.0029 and 0.0017 in the global sample.
+Similarly, for the same time windows, the post-news coefficients for negative news sentiment are -0.0089 and -0.0066 in Japan, compared with -0.0046 and -0.0025 in the global sample.
+Second, the post-news coefficient for positive news sentiment is smaller than its pre-news coefficient (Panel (A) of Figure 9), although the post-news coefficient is always larger in size than the pre-news coefficient in the global sample.
+This finding implies that positive information about a firm may be incorporated into prices prior to disclosure, for example suggestions of financial companies, SNS, and word of mouth, and that such information is associated with higher stock prices of the firm’s suppliers.
+However, once the positive information is disclosed in a news article, the magnitude of the coefficients for suppliers is smaller than that in the pre-news period.
+In other words, for Japanese firms, the difference between the post- and pre-news effects on suppliers is negative, indicating that the pre-news effect exceeds the post-news effect.
+In contrast to the results for positive news, the response to negative news in Japan exhibits the same qualitative pattern as that observed in the global sample.
+Specifically, for negative news, the post-news effect is larger than the pre-news effect for Japanese firms, as is also the case for firms worldwide (Panel (B) of Figure 9 and Panel (B) of Figure 6).
+Within this common pattern, however, there are differences across time windows.
+In particular, for Japanese firms, the pre-news effect of negative news for the one-day time window is smaller in absolute value than that for longer windows of two to five days, whereas the global results show an increase in the absolute value of the pre-news effect as the window size decreases.
+We further examine coefficients for news sentiment on client firms and present the results in Figure 10.
+Detailed regression results are reported in SI Tables 12 and 13.
+Compared with the corresponding results for firms across the world shown in Figure 7, we find that the same three differences observed for suppliers also apply to clients.
+First, the average magnitude of the effects is generally larger for Japanese firms than for firms in the global sample.
+Second, for positive news, the difference between the post- and pre-news effects on clients is negative in Japan.
+Third, for negative news, the pre-news effect for the one-day time window is smaller than that for longer time windows.
+5 Discussion
 In summary, we find that the way positive and negative information about firms affects stock prices of their suppliers and clients before and after the disclosure of the information by news media is different between firms in Japan and those in other countries.
-Specifically, we find that positive news propagates more before its disclosure than after in the Japanese data.
-Although we cannot provide evidence for any mechanism behind the differences, one possible reason is the difference in the quality of supply chain information.
-The supply chain data used for firms worldwide in this study relies on financial statements, whereas the data for Japan is based on a questionnaire-based survey.
-Although the former is automatically determined by criteria based on sales volume, the latter identifies important trade partners, which are not determined only by sales volume.
-Since investors who can receive information about firms through private channels are more knowledgeable about supply chains and recognize the diffusion of performance through supply chains more accurately than other investors who rely on news media, once the knowledgeable investors receive positive information about a firm before its disclosure by news media, they invest in the firm’s suppliers and clients.
-This results in a large increase in stock prices of the suppliers and clients before the disclosure of the news.
-However, other investors who rely on news media and do not understand the role of supply chains react less strongly to positive news after its disclosure.
-Another possible reason for this is the distinct nature of Japan’s supply chains, known as keiretsu.
-Links between suppliers and clients in these relationships are often stronger than those in other countries, as they are frequently associated with shareholding relationships, information exchanges, technical assistance, and research collaboration [67, 68].
-These strong relationships may explain why positive news about firms may have a greater impact on their suppliers and clients.
-Because investors recognize the strength of these ties and anticipate the transmission of performance through supply chains, news about firms tends to have a greater impact on the stock prices of their suppliers and clients in Japan compared to other countries.
-However, these two reasons cannot explain why negative news does not propagate before disclosure, which should be investigated in future research.
+Additionally, we find that stronger pre-news stock price responses are observed for positive news in the Japanese sample.
+It is important to emphasize that these results do not imply that investors possess detailed knowledge of firm-level supply chain structures.
+Rather, the observed stock price responses are consistent with market pricing that reflects information correlated with supply chain relationships.
+Several factors may be correlated with the observed stock price responses along supply chain relationships.
+The present analysis does not identify the mechanisms underlying these correlations.
+It is also important to distinguish between realized economic impacts and market-based expectations reflected in stock prices.
+The present study focuses on stock price responses to news, which capture investors’ expectations about potential spillovers along supply chain relationships rather than confirmed changes in production, orders, or real economic outcomes.
+Whether such expectations are eventually realized in the form of actual production disruptions or demand reductions is beyond the scope of this analysis.
+In this context, we clarify the specific dimension of market responses examined in this study.
+The focus of this study is on stock price responses, which reflect changes in market valuation.
+While other market indicators such as trading volume or volatility capture important aspects of market activity and uncertainty, they represent different dimensions of market behavior and are not direct measures of valuation effects.
+Accordingly, the present analysis concentrates on price-based responses to quantify how news-related information propagates through supply chain networks.
+Analyses of news spillovers in trading volume or volatility would require a different empirical design and are therefore not pursued in this study.
+With this focus on price-based responses in mind, we turn to the interpretation of differences between the global and Japanese samples.
+These differences between the global and Japanese samples do not imply that different databases mechanically generate different results.
+Rather, the results consistently show that stock price responses propagate through supply chain relationships in both samples, while the timing and relative magnitude of such propagation differ across institutional contexts.
+The use of different supply chain databases reflects these institutional contexts rather than causing the observed differences.
+The FactSet database captures major supply chain links disclosed in financial statements, whereas the TSR database identifies economically important trading partners through firm-level surveys.
+These databases therefore provide complementary views of supply chain relationships, rather than alternative measurements of the same underlying network.
+Although the FactSet database is mainly determined by criteria based on sales volume, the TSR database identifies important trade partners, which are not determined only by sales volume.
+One possible interpretation is that the observed differences are consistent with heterogeneous timing and magnitude of market responses across institutional contexts.
+The present analysis does not establish the mechanisms generating these patterns.
+Several institutional and historical factors may potentially contribute to the observed differences between Japan and other countries.
+The observed differences arise in an institutional context in which supply chain relationships in Japan are often discussed using the term keiretsu [3, 50].
+The present analysis does not identify whether, or how, such institutional features contribute to the observed patterns.
+Other factors, such as historical shocks to domestic supply chains or differences between domestic and global supply chain structures, may also play a role.
+Identifying the relative importance of these mechanisms is beyond the scope of this study and remains an important direction for future research.
 6 Conclusion
-This study examines how positive and negative news about firms affects the change rate of their stock prices and, moreover, how it affects the change rate of stock prices of the firms’ suppliers and clients, using a large sample of publicly listed firms across the world and another of Japanese listed firms.
-The level of positiveness and negativeness of each news article is determined by FinBERT, a natural language processing model fine-tuned specifically for financial information.
-Supply chains of firms across the world are identified mostly by financial statements and supplemented by other information on the website and news medias, while those of Japanese firms are taken from large-scale firm-level surveys.
-We find that positive news increases the change rate of stock prices of firms mentioned in the news before its disclosure, most likely because of diffusion of information through private channels, such as SNS and word of mouth.
-Positive news also raises stock prices of the firms’ suppliers and clients before and after its disclosure, confirming propagation of market values through supply chains.
-In addition, we generally find a larger post-news effect on stock prices of the mentioned firms and their suppliers and clients than the pre-news effect.
-The positive difference between the post- and pre-news effects can be considered as the net effect of the disclosure of positive news, controlling for information diffusion through private channels.
-However, the post-news effect on suppliers and clients in Japan can be smaller than the pre-news effect, a result opposite to those from firms across the world.
-This notable result is possibly because supply chain links of Japanese firms are stronger than global supply chains while such knowledge is restricted to selected investors.
+We find that news sentiment, both positive and negative, is associated with systematically higher and lower change rates of stock prices of firms mentioned in the news in the pre-news period, which may be consistent with a combination of anticipatory market behavior, information correlated with forthcoming news, and selection in news coverage.
+Such news sentiment is also associated with stock price responses of the firms’ suppliers and clients in both the pre-news and post-news periods, consistent with systematic valuation comovements along supply-chain relationships.
+In addition, we generally find that post-news effects on stock prices of the mentioned firms and their suppliers and clients are larger than pre-news effects.
+The difference between post- and pre-news effects summarizes differences in stock price changes between the post- and pre-disclosure periods.
+However, for Japanese firms, post-news effects on suppliers and clients can be smaller than pre-news effects, in contrast to the pattern observed for firms across the world.
 The important opening question not addressed by this study is the significant heterogeneity of supply chain relationships, which can lead to varying impacts of news sentiment.
@@
 Data availability
-The data that support the findings of this study are available from FactSet Research Systems Inc.
-and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.
-but restrictions apply to the availability of these data, which were used under license for the current study, and so are not publicly available.
-Data are however available from the authors upon reasonable request and with permission of FactSet Research Systems Inc.
-and Tokyo Shoko Research, Ltd, and London Stock Exchange Group Data & Analytics.
+The data used in this study are obtained from commercial databases, including FactSet Research Systems Inc., Tokyo Shoko Research, Ltd., and London Stock Exchange Group Data & Analytics (Refinitiv).
+These data are subject to licensing restrictions and cannot be made publicly available by the authors.
+Researchers with access to the same commercial data sources under appropriate licenses can obtain the data directly from the respective providers and reproduce the analyses conducted in this study.
+The authors do not have special access privileges to these data beyond standard institutional licenses.
 Acknowledgements
 The authors thank Wataru Souma and Takayuki Yumoto for valuable discussions and Yuito Nishi for his technical contribution.
+Funding information
+This work was supported by JSPS KAKENHI Grant Numbers JP22K18533, JP23K20626, JP23K25520, JP25K01454, JST PRESTO Grant Number JPMJPR21R2, and the Asahi Glass Foundation.
 Author contributions statement
-H.I conceived the study. H.I and Y.T conducted the analyses and wrote and reviewed the manuscript.
+H.I conceived the study. H.I and Y.T conducted the analyses, and wrote and reviewed the manuscript.